# thanks: These two authors contribute equallythanks: These two authors contribute equally

### v1

@@ Abstract
For a given parity check matrix $H_{Z}=\begin{pmatrix}H_{A}&H_{B}\\ \end{pmatrix}$, for which $\text{rank}(H_{A})=r_{A}$, $\text{rank}(H_{B})=r_{B}$ and $\text{rank}(H_{Z})=r_{H}$, it can be proved that (see Appendix E for details)
@@
This provides an efficient scheme for evaluating the entanglement entropy, since the rank of a matrix can be calculated efficiently. The entanglement entropy for other logical states can be evaluated similarly by appending rows representing logical Pauli $Z$ operators to the parity check matrix $H_{Z}$ (see Appendix F for details).
Figure 2: Entanglement entropy of BB codes and toric codes. (a) Scaling behavior of entanglement entropy for subsystems with increasing sizes. The size of the subsystem is chosen to be less than half of the entire system. (b) Averaged entanglement entropy for randomly selecting subsystems. (c) Derivative of the discrepancy of entropy $I_{A}$ for BB codes. The inset shows the same quantity for toric codes.
@@
Entanglement entropy for qLDPC codes. We apply the graph-based method to calculate the entanglement entropy for a family of qLDPC codes called bivariate bicycle (BB) codes Bravyi et al. (2024b). They are quantum codes of CSS type for which each stabilizer generator acts non-trivially on six qubits and each qubit participates in six stabilizer generators. The BB codes are similar to the two-dimensional toric codes except that they are not geometrically local. We show that the entanglement properties of the BB codes differ significantly from those of the two-dimensional toric codes.
A subsystem with a convex disk shape in toric codes encloses nearly the maximum number of stabilizer generators. We select a series of subsystems in BB codes with a similar property (see Appendix G for details).
@@
The von Neumann entropy of the defined subsystems provides insight into their scaling behavior. The entanglement entropy for a toric code with $d=20$ and a BB code [[756,16,34]][[756,16$\leq$ 34]][ [ 756 , 16 , ≤ 34 ] ] (see Appendix H for code construction) is shown in Fig. 2 (a). The entanglement entropy for the toric code is approximately proportional to $\sqrt{n_{A}}$, consistent with an area law that reflects the localized nature of its entanglement. In contrast, for the BB code, the entanglement entropy scales as $n_{A}^{\gamma}$ with $\gamma\approx 0.81$, showing a faster increase in the entanglement entropy and revealing the presence of more delocalized and long-range entanglement.
Another way to choose a subsystem $A$ is to randomly select a set of qubits from the code. This is in some sense equivalent to fixing a subsystem and averaging over a set of different quantum states. The average entanglement entropy for both toric codes and BB codes is shown in Fig. 2 (b). It is evident that the average entanglement entropy $\bar{S}_{A}$ exhibits a linear increase with respect to the size of the subsystem, $n_{A}$, in the regime where $n_{A}$ is small, suggesting a volume law. It deviates from the volume law when $n_{A}$ approaches half of the total system size and the deviation reaches a maximum at $n_{A}=N/2$. It starts to decrease when $n_{A}>N/2$, and exhibits a linear decrease as $n_{A}$ approaches $N$.
It may seem counterintuitive that the average entanglement entropy of small random subsystems follows a volume law.
This is due to the unique characteristics of the stabilizer codes. When $n_{A}$ is smaller than the number of qubits on which a stabilizer generator acts non-trivially, the state of the subsystem is maximally mixed. As a result, the entanglement entropy contributed by this subsystem is $n_{A}$. When $n_{A}$ exceeds the number of qubits on which a stabilizer generator acts non-trivially, and some of these qubits form one or more stabilizer generators (cycles in the language of graph theory), the entanglement entropy deviates from $n_{A}$. We define the discrepancy between $n_{A}$ and $\bar{S}_{A}$ as a quantitative measure that characterizes the deviation, $I_{A}=n_{A}-\bar{S}_{A}(n_{A})$.
The entropy discrepancy $I_{A}$ depends on the number of independent stabilizer generators that act exclusively within subsystem $A$.
When a randomly selected subsystem is small compared to the entire system, the selected quibts are likely to be disconnected from each other, forming many isolated qubit “islands”. The entanglement entropy of each qubit island is typically equal to its size, with high probability, leading to the observed volume law. The entanglement entropy increases linearly with $n_{A}$ as long as the size of the isolated qubit island remains sufficiently small. Consequently, adding more qubits to subsystem $A$ does not violate the volume law.
As the size of a randomly selected subsystem grows so that a qubit island encompasses one or more stabilizer generators, the average entanglement entropy starts to deviate from the volume law. Beyond this point, adding more qubits to the subsystem increases its entropy discrepancy, and as the subsystem grows further, the rate of change of the entropy discrepancy approaches a constant.
We calculate the rate of change of the entropy discrepancy, ${\rm d}I_{A}/{\rm d}n_{A}$, for both BB codes and toric codes, revealing a significant difference between the two, as shown in Fig. 2 (c). A sharp transition from a phase with a vanishing rate to one with a constant rate is observed for the BB codes, with the transition becoming sharper as the size of the BB codes increases. In contrast, for the toric codes, the transition is much smoother and the steepness of the transition is independent of the size of the toric codes. The difference can be qualitatively understood as follows. The stabilizer generators for the toric codes act non-trivially on four qubits, whereas those for the BB codes act non-trivially on six qubits. As a result, the isolated qubit islands for the BB codes must grow larger than those for the toric codes to break the volume law.
The sharp transition for the rate of change of the entropy discrepancy for the BB codes can be qualitatively understood using concepts from percolation theory Broadbent and Hammersley (1957). A random subsystem of size $n_{A}$ is obtained by randomly selecting $n_{A}$ qubits from a total of $N$ qubits, with each quibt being selected with probability $p_{s}=n_{A}/N$ and discarded with probability $1-p_{s}$. When $p_{s}$ is small, the qubits in subsystem $A$ form many disconnected small qubit islands. However, when $p_{s}$ is sufficiently large, the qubits in subsystem $A$ merge into a few connected, large qubit islands. According to percolation theory, there exists a percolation threshold below which the quibts remain disconnected and above which quibts rapidly become connected. The connectivity of qubits in subsystem $A$ is related to the entropy discrepancy. This implies a close relationship between the phase transition in percolation theory and the observed sharp transition for the rate of change of the entropy discrepancy.
@@
Conclusions. We developed a graph-based method to explore the density matrix and the entanglement entropy of CSS-type quantum error correcting codes. The entanglement entropy has a simple graph-theoretical interpretation: it is the cyclomatic number of the cycle space formed by joining the spanning trees of the subsystem and its complementary subsystem. Its analytic expression related to the matrix rank inspires an efficient scheme to calculate the entanglement entropy for complicated quantum codes, e.g., qLDPC codes. We demonstrated the method by calculating and comparing the entanglement entropy for the toric codes and the BB codes. The graph-based method does not depend on the specific geometry and topology of the quantum systems and can be generalized to explore systems with higher spins and irregular structures, as well as a broader class of stabilizer codes.
@@ Appendix A Edge space, cycle space and symmetric difference
One can extract a vector space from a graph. The process involves extracting the edge set from a graph, which subsequently allows for the construction of a corresponding vector space based on that set. In graph theory, this space is referred to as the edge space. Before constructing the vector space, it is essential to introduce the concept of symmetric difference between two sets. Given two sets $A$ and $B$, the symmetric difference between $A$ and $B$ is the set of elements that are in either $A$ or $B$, but not in their intersection. This is denoted as $A\Delta B$ and can be expressed mathematically as
@@
The cycle space of a graph $G$ consists of all cycles in the graph, denoted as $c(G)$. It is a subspace of the edge space. The vector addition of cycle space also amounts to the symmetric difference. The basis of the cycle space is the minimal set of cycles which can span all the cycles in the graph. The dimension of the cycle space is defined as the number of elements in the basis, which is called cyclomatic number.
@@ Appendix C Some examples in toric codes
Figure 4: Illustration of subsystems in the toric code, and their spanning trees and joint cycles. These subsystems are the qubits on the red edges. (a) The spanning tree of subsystem $A$ is obtained by removing one of the $d$ qubits (red cross) along the non-contractible qubit chain. The spanning tree of subsystem $B$ is obtained by removing appropriate qubits (black cross) in subsystem $B$. The joint cycles are indicated by the orange faces. (b) The spanning forest of subsystem $A$ consists of the qubits on the red edges. The spanning tree of subsystem $B$ is obtained by removing appropriate qubits (black cross) in subsystem $B$. The qubits of $B$ on the blue edges form non-contractible joint cycles with the edges in the spanning forest of $A$. (c) The red edges, with the exception of the one that has been removed (red cross), constitute the spanning forest of subsystem $A$. By removing the edges indicated by the black cross, one can obtain the spanning tree of subsystem $B$. The spanning forest of $A$ and the spanning tree of $B$ share the stabilizer operators that are indicated by one pink face and $d-2$ orange faces. The red edges and the blue edges also form non-contractible joint cycles. (d) The spanning forest of subsystem $A$ and that of subsystem $B$ form a joint cycle space. Two spanning forests share the stabilizer operators which are indicated by the orange faces.
We calculate the entanglement entropy for several types of subsystem for the logical state $|00\rangle_{L}$, including a single qubit, two qubits, the qubit chain, the vertical qubit ladder, the cross and all vertical qubits. Some of these subsystems are shown in Fig. 4.
@@ C.3 The qubit chain
Consider a $d\times d$ square lattice and a non-contractible qubit chain in the torus, as shown by the red edges in Fig. 4 (a). The spanning tree of subsystem $A$ and that of subsystem $B$ form a joint cycle space with dimension $d-1$. Therefore, the entanglement entropy for the non-contractible qubit chain is $d-1$.
@@ C.5 The cross
The qubits in subsystem $A$ form a cross, as shown by the red edges in Fig. 4 (c). The spanning forest of subsystem $A$ and the spanning tree of subsystem $B$ form a joint cycle space. A basis of the joint cycle space consists of $d-1$ contractible cycles, represented by a pink face and orange faces, along with $d$ non-contractible cycles.The cyclomatic number of the cycle space is given by $12d-$ 1, which directly corresponds to the entanglement entropy of the system. Hence, the entanglement entropy is $12d-$ 1.
@@ C.6 All vertical qubits
The subsystem $A$ contains all vertical qubits in the torus, as shown by the red edges in Fig. 4 (d). By removing all the vertical edges in the first row, one can obtain the spanning forest of $A$. Similarly, the spanning forest of $B$ can be obtained by removing all horizontal edges from the first column. The spanning forest of $A$ and that of $B$ form a cycle space with dimension $(d-1)^{2}$, as shown in Fig. 4 (d). The entanglement entropy is $(d-1)^{2}$.
@@ Appendix D Transformation of parity check matrix
Figure 5: The process for transforming the parity check matrix. Prior to step 1, appropriate row and column permutations are applied to reorganize the parity check matrix. Step 1 involves deleting the cycles within subsystem $A$ that do not intersect with the boundary of subsystem $B$, as well as the cycles within subsystem $B$ that do not intersect with the boundary of subsystem $A$. Step 2 focuses on eliminating the cycles that lie on the boundaries of both subsystems $A$ and $B$. Step 3 entails transforming the parity check matrix into its final form through Gaussian elimination.
We provide a detailed process of transforming the parity check matrix $H_{Z}$, as summarized in Fig. 5. Once subsystems $A$ and $B$ are defined, the parity check matrix $H_{Z}$ of a CSS code can be partitioned into $(H_{A}~{}H_{B})$ through appropriate column permutations, where $H_{A}$ and $H_{B}$ are submatrices whose columns correspond to qubits in subsystems $A$ and $B$, respectively. By further appropriate row permutations, the parity check matrix can be transformed into
@@ Appendix F Entanglement entropy for logical states
|$\ket{\psi_{L}}=\frac{1}{2^{k/2}}\sum\limits_{C\in\{0,1\}^{k}}\ket{C}_{L},$ (36)
where |$\ket{C}_{L}$ represented a logical computational basis state, and $k$ denotes the number of logical qubits. In fact, our method can also be used to evaluate the entanglement entropy for an equal superposition of a subset of logical computational bases. Denote the logical Pauli $Z$ operators as $\bar{Z}_{1},\bar{Z}_{2},\cdots,\bar{Z}_{k}$, and the corresponding values of logical qubits as $\bar{z}_{1},\bar{z}_{2},\cdots,\bar{z}_{k}$, with $\bar{z}_{i}=\{0,1\}$. If one of the logical qubits takes a specific value, say $\bar{z}_{1}=0$, then the encoded state is subject to an additional constraint. By appending a row representing the logical operator $\bar{Z}_{1}$ to the parity check matrix $H_{Z}$, we can evaluate the entanglement entropy for the state
|$\ket{\psi_{L}^{(1)}}=\frac{1}{2^{(k-1)/2}}\sum_{\bar{z}_{2}}\sum_{\bar{z}_{3}}\cdots\sum_{\bar{z}_{k}}\ket{0,\bar{z}_{2},\cdots,\bar{z}_{k}}_{L}.$ (37)
@@
If the tensor product (in the code subspace) of two logical Pauli $Z$ operators, say $\bar{Z}_{1}\otimes\bar{Z}_{2}$, takes a specific eigenvalue, say $+1$, then the encoded state is subject to an additional constraint. By appending a row representing the logical operator $\bar{Z}_{1}\otimes\bar{Z}_{2}$ to the parity matrix $H_{Z}$, we can evaluate the entanglement entropy for the state
|$\ket{\psi_{L}^{(12)}}$ $=$ $\frac{1}{2^{(k-1)/2}}\bigg{\{}\sum_{\bar{z}_{3}}\sum_{\bar{z}_{4}}\cdots\sum_{\bar{z}_{k}}\ket{0,0,\bar{z}_{3},\cdots,\bar{z}_{k}}_{L}$ (38)
$+\sum_{\bar{z}_{3}}\sum_{\bar{z}_{4}}\cdots\sum_{\bar{z}_{k}}\ket{1,1,\bar{z}_{3},\cdots,\bar{z}_{k}}_{L}\bigg{\}}.$
@@
In the special case where every logical Pauli $Z$ operator takes a specific eigenvalue, we actually evaluate the entanglement entropy for a logical computational basis state |$\ket{C}_{L}$. It can be proved that the entanglement entropy for any logical basis state is the same as that of the logical basis state |$\ket{00\cdots 0}_{L}$, for which the eigenvalues of all logical Pauli $Z$ operators are equal to one Kargarian (2008). This equivalence arises from the fact that any logical basis state |$\ket{C}_{L}$ can be transformed into |$\ket{00\cdots 0}_{L}$ by applying some logical Pauli $X$ operators to flip the corresponding logical qubits from 00 to $1$, and any logical Pauli $X$ operator is a tensor product of single-qubit Pauli $X$ operators.

### v3

@@ Abstract
For a given parity check matrix $H_{Z}=\begin{pmatrix}H_{A}&H_{B}\\ \end{pmatrix}$, for which $\text{rank}(H_{A})=r_{A}$, $\text{rank}(H_{B})=r_{B}$ and $\text{rank}(H_{Z})=r_{H}$, it can be proved that SI (2025)
@@
This provides an efficient scheme for evaluating the entanglement entropy, since the rank of a matrix can be calculated efficiently. The entanglement entropy for other logical states can be evaluated similarly by appending rows representing logical Pauli $Z$ operators to the parity check matrix $H_{Z}$ SI (2025).
@@
Entanglement entropy for qLDPC codes. We apply the graph-based method to calculate the entanglement entropy for two families of qLDPC codes: bivariate bicycle (BB) codes Bravyi et al. (2024) and quasi-cyclic (QC) codes Hagiwara and Imai (2007), both of which are quantum codes of CSS type. In BB codes, each stabilizer generator acts non-trivially on six qubits and each qubit participates in six stabilizer generators; while QC codes exhibit variable stabilizer weights. Neither BB codes nor QC codes are geometrically local. We show that the entanglement properties of the BB codes and QC codes differ significantly from those of the two-dimensional toric codes SI (2025).
Figure 2: Scaling of entanglement entropy for toric codes, bivariate bicycle codes and quasi-cyclic codes. The size of the subsystem is chosen to be less than half of the entire system.
@@
The entanglement entropy for a toric code with code distance $d=20$, a BB code $756,16,\leq 34]]$ and a QC code $710,8,\leq 96]]$ is shown in Fig. 2. In this calculation, the subsystem $A$ is selected in a specific way such that it remains a single connected component as its size increases SI (2025). The entanglement entropy for the toric code is approximately proportional to $\sqrt{n_{A}}$, consistent with an area law that reflects the localized nature of its entanglement. In contrast, for the BB and QC codes, the entanglement entropy scales as $n_{A}^{\gamma}$ with $\gamma\approx 0.81$ and $0.95$, respectively, showing a faster growth in the entanglement entropy and revealing the presence of more delocalized and long-range entanglement.
We then randomly select subsystems $A$ and calculate their average von Neumann entropy $\bar{S}_{A}$, also called the average entanglement entropy. The results for toric codes and BB codes are shown in Fig. 3 (a). It can be seen that the average entanglement entropy for sufficiently small random subsystems follows a volume law. This is due to the unique characteristics of the stabilizer codes. When a randomly selected subsystem is small compared to the entire system, the selected quibts are likely to be disconnected from each other, forming many isolated qubit “islands”. The state of the qubit island is maximally mixed, and its entanglement entropy is typically equal to its size, leading to the observed volume law. As the size of a randomly selected subsystem grows such that a qubit island encompasses one or more stabilizer generators (cycles in the language of graph theory), the average entanglement entropy starts to deviate from the volume law. We define the discrepancy between subsystem size $n_{A}$ and $\bar{S}_{A}$ as a quantitative measure to characterize the deviation, $I_{A}=n_{A}-\bar{S}_{A}(n_{A})$. The entropy discrepancy $I_{A}$ depends on the number of independent stabilizer generators that act exclusively within subsystem $A$.
Figure 3: Entanglement entropy for qLDPC codes and toric codes. (a) Average entanglement entropy for randomly selecting subsystems. (b) Derivative of the discrepancy of entropy $I_{A}$ for $B$ codes. The inset shows the same quantity for toric codes. (c) Derivative of $I_{A}$ for QC codes with stabilizer weight 6. (d) Derivative of $I_{A}$ for QC codes with stabilizer weight 8.
We calculate the rate of change of the entropy discrepancy, ${\rm d}I_{A}/{\rm d}n_{A}$, for both toric codes and qLDPC codes, revealing a significant difference between the two, as shown in Figs. 3 (b)-(d). A sharp transition from a phase with a vanishing rate to one with a constant rate is observed for the qLDPC codes, with the transition becoming sharper as the size of the codes increases. In contrast, for the toric codes, the transition is much smoother and the steepness of the transition is independent of the size of the toric codes. The difference can be qualitatively attributed to variations in connectivity. Toric codes are defined on two-dimensional surfaces and exhibit strictly local interactions. As a result, the connectivity between subsystems, which is directly related to the entanglement entropy (see Eq. (5)), remains independent of the code size. In contrast, the high stabilizer weights and non-local interactions in BB and QC codes enhance connectivity between subsystems, leading to distinct entanglement properties. This is also evident, by comparing Figs. 3 (c) and (d), from the observation that QC codes with higher stabilizer weights display a sharper transition.
The sharp transition in the rate of change of the entropy discrepancy in BB and QC codes reflects a distinct mechanism compared to toric codes. When subsystem $A$ is much smaller than half of the entire code and subsystem $B$ is a single connected component, transferring one qubit from $B$ to $A$ increases the entropy discrepancy only when a cycle or an $X$-type stabilizer is formed in $A$ SI (2025), the likelihood of which depends on both the stabilizer weight and the connectivity of the code. A larger stabilizer weight implies that subsystem $A$ must grow to a greater size before a cycle or an $X$-type stabilizer can be formed via adding a qubit, leading to a sharper transition. Due to the local nature of the toric code, its connectivity remains invariant as the code size increases, giving rise to ${dI_{A}}/{dn_{A}}$ that is independent of the code size. In contrast, for qLDPC codes, the connectivity evolves with the code size such that qubits participating in a stabilizer become more spatially separated (if embedded in a two-dimensional plane), rendering the formation of cycles or $X$-type stabilizers in $A$ increasingly difficult. This behavior serves as a clear empirical signature of the fundamental difference in entanglement properties between locally and nonlocally constrained quantum codes.
@@
Conclusions. We developed a graph-based method to explore the density matrix and the entanglement entropy of CSS-type quantum error correcting codes. The entanglement entropy has a simple graph-theoretical interpretation: it is the cyclomatic number of the cycle space formed by joining the spanning trees of the subsystem and its complementary subsystem. Its analytic expression related to the matrix rank inspires an efficient scheme to calculate the entanglement entropy for complicated quantum codes, e.g., qLDPC codes. We demonstrated the method by calculating and comparing the entanglement entropy for the toric codes and two families of qLDPC codes: the BB and QC codes, revealing significantly distinct entanglement properties between these two sets of quantum codes. These entanglement properties could be connected to the information protection capacity and decoding performance of CSS codes, which warrant further investigation in future work. The graph-based method does not depend on the specific geometry and topology of the quantum systems and can be generalized to explore systems with higher spins and irregular structures, as well as a broader class of stabilizer codes.
@@ Appendix A Edge space, cycle space and symmetric difference
One can extract a vector space from a graph. The process involves extracting the edge set from a graph, which subsequently allows for the construction of a corresponding vector space based on that set. In graph theory, this space is referred to as the edge space Diestel (2024). Before constructing the vector space, it is essential to introduce the concept of symmetric difference between two sets. Given two sets $A$ and $B$, the symmetric difference between $A$ and BB is the set of elements that are in either $A$ or $B$, but not in their intersection. This is denoted as $A\Delta B$ and can be expressed mathematically as
@@
The cycle space of a graph $G$ consists of all cycles in the graph, denoted as $c(G)$. It is a subspace of the edge space. The vector addition of cycle space also amounts to the symmetric difference. The basis of the cycle space is the minimal set of cycles which can span all the cycles in the graph, with the elements of this set being defined as independent cycles. The dimension of the cycle space is defined as the number of elements in the basis, which is called cyclomatic number.
@@ Appendix C Some examples in toric codes
Figure 5: Illustration of subsystems in the toric code, and their spanning trees and joint cycles. These subsystems are the qubits on the red edges. (a) The spanning tree of subsystem $A$ is obtained by removing one of the $d$ qubits (red cross) along the non-contractible qubit chain. The spanning tree of subsystem $B$ is obtained by removing appropriate qubits (black cross) in subsystem $B$. The joint cycles are indicated by the orange faces. (b) The spanning forest of subsystem $A$ consists of the qubits on the red edges. The spanning tree of subsystem $B$ is obtained by removing appropriate qubits (black cross) in subsystem $B$. The qubits of BB on the blue edges form non-contractible joint cycles with the edges in the spanning forest of $A$. (c) The red edges, with the exception of the one that has been removed (red cross), constitute the spanning forest of subsystem $A$. By removing the edges indicated by the black cross, one can obtain the spanning tree of subsystem $B$. The spanning forest of $A$ and the spanning tree of BB share the stabilizer operators that are indicated by one pink face and d2d-2 orange faces. The red edges and the blue edges also form non-contractible joint cycles. (d) The spanning forest of subsystem $A$ and that of subsystem $B$ form a joint cycle space. Two spanning forests share the stabilizer operators which are indicated by the orange faces.
We calculate the entanglement entropy for several types of subsystem for the logical state $|00\rangle_{L}$, including a single qubit, two qubits, the qubit chain, the vertical qubit ladder, the cross and all vertical qubits Hamma et al. (2005b). Some of these subsystems are shown in Fig. 5.
@@ C.3 The qubit chain
Consider a $d\times d$ square lattice and a non-contractible qubit chain in the torus, as shown by the red edges in Fig. 5 (a). The spanning tree of subsystem $A$ and that of subsystem $B$ form a joint cycle space with dimension d1d-1. Therefore, the entanglement entropy for the non-contractible qubit chain is d1d-1.
@@ C.5 The cross
The qubits in subsystem $A$ form a cross, as shown by the red edges in Fig. 5 (c). The spanning forest of subsystem $A$ and the spanning tree of subsystem $B$ form a joint cycle space. A basis of the joint cycle space consists of d1d-1 contractible cycles, represented by a pink face and orange faces, along with $d$ non-contractible cycles.The cyclomatic number of the cycle space is given by 2d12d-1, which directly corresponds to the entanglement entropy of the system. Hence, the entanglement entropy is 2d12d-1.
@@ C.6 All vertical qubits
The subsystem $A$ contains all vertical qubits in the torus, as shown by the red edges in Fig. 5 (d). By removing all the vertical edges in the first row, one can obtain the spanning forest of $A$. Similarly, the spanning forest of BB can be obtained by removing all horizontal edges from the first column. The spanning forest of $A$ and that of BB form a cycle space with dimension $(d-1)^{2}$, as shown in Fig. 5 (d). The entanglement entropy is $(d-1)^{2}$.
@@ Appendix D Transformation of parity check matrix
Figure 6: The process for transforming the parity check matrix. Prior to step 1, appropriate row and column permutations are applied to reorganize the parity check matrix. Step 1 involves deleting the cycles within subsystem AA that do not intersect with the boundary of subsystem $B$, as well as the cycles within subsystem $B$ that do not intersect with the boundary of subsystem $A$. Step 2 focuses on eliminating the cycles that lie on the boundaries of both subsystems $A$ and $B$. Step 3 entails transforming the parity check matrix into its final form through Gaussian elimination.
We provide a detailed process of transforming the parity check matrix $H_{Z}$, as summarized in Fig. 6. Once subsystems $A$ and BB are defined, the parity check matrix $H_{Z}$ of a CSS code can be partitioned into $(H_{A}~H_{B})$ through appropriate column permutations, where $H_{A}$ and $H_{B}$ are submatrices whose columns correspond to qubits in subsystems $A$ and $B$, respectively. By further appropriate row permutations, the parity check matrix can be transformed into
@@ Appendix F Entanglement entropy for logical states
[…]
@@
If the tensor product (in the code subspace) of two logical Pauli $Z$ operators, say $\bar{Z}_{1}\otimes\bar{Z}_{2}$, takes a specific eigenvalue, say +1+1, then the encoded state is subject to an additional constraint. By appending a row representing the logical operator $\bar{Z}_{1}\otimes\bar{Z}_{2}$ to the parity matrix $H_{Z}$, we can evaluate the entanglement entropy for the state
[…]
@@
[…]